OpenStudy (anonymous):

If F(x)=x^2+1 and G(x)= 2x+3 write the composition for: G(F(x))_______, F(G(x))_______

OpenStudy (anonymous):

G(F(X))=G(X^2+1)=2(X^2+1)+3=?

OpenStudy (anonymous):

lol im not sure

OpenStudy (anonymous):

=2x^2+5

OpenStudy (anonymous):

for which one?

OpenStudy (anonymous):

similarly F(G(X))=F(2X+3)=(2X+3)^2+1 solve it.